OpenStudy (anonymous):

Basis of ( will provide vector) in R^3

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ 3\\4\end{matrix}\right),\left(\begin{matrix}4 \\ 3\\8\end{matrix}\right)\]